# EasyOcc: 3D Pseudo-Label Supervision for Fully Self-Supervised Semantic Occupancy Prediction Models

### v3

@@ V-E5 Lovàsz Softmax Loss
Following the previous experiment, we retrain the model with the Lovász-Softmax loss that excludes the empty index from the loss computation. The results are presented in Table VIII.
TABLE VIII: Ablation: Lovász-Softmax Empty Index: Models are trained for 12 epochs to ensure manageable training time. The best-performing configuration is highlighted in bold.
@@
13.18
13.55
Excluding the empty class from the loss computation results in a 0.37 mIoU improvement. Only the previous experiments in the ablation section were conducted with the inclusion of the empty label in the Lovász loss. However, the integrity of the comparative results is maintained, as all models were subject to the same conditions and limitations.
@@ V-F Qualitative Results
In this section, we present a qualitative analysis in two parts. First, we evaluate the quality of predicted semantic voxels for all models compared to the ground-truth labels. Following this, we examine the voxel predictions from each of the six camera views, comparing GaussianOcc to its variant trained with our proposed pseudo-loss to examine the benefits of our 3D pseudo-labels. TT-OccCamera, GaussianFlowOcc, and AutoOcc are omitted because they are not currently open-source.
@@ V-F1 Voxel Analysis
Figure 8: Voxel Qualitative Analysis: Visual analysis of models evaluated in Table III. This analysis is conducted on nuScenes scene-783, sample token e67f3e81225f426f8e1743af45487762. Boxes are overlaid to highlight specific areas for qualitative analysis.
We begin the qualitative analysis by inspecting Figure 8, which shows semantic voxel visualizations for each model in the SOTA comparison in Table IV. These visualizations provide insight into the strengths and limitations of each model, starting with a comparison of the original models and their counterparts augmented with our proposed pseudo-loss.
SelfOcc performs well in road segmentation but misses vehicles in the back camera view. With the implementation of pseudo-loss, the model correctly identifies these vehicles. However, with the inclusion of pseudo-loss, we observe a notable increase in misclassifications as the barrier class, particularly in the back-left camera view—an issue not present in the original model. This may be attributed to the presence of a metal fence and a construction road sign in the corresponding image.
OccNeRF predicts more elements than SelfOcc, including a car and a bus in the rear view, but introduces noticeable object duplication, which results in heavy penalisation when evaluating without the camera mask. Pseudo-loss reduces this duplication and noticeably improves vegetation identification in the front camera view. A similar improvement is seen in GaussianOcc, where pseudo-loss corrects mispredictions, such as classifying road as wall in the rear view.
Comparing our model, EasyOcc, to GaussTR, key differences emerge. GaussTR fails to predict vegetation in the front camera view, whereas EasyOcc does. Additionally, the overhang of a building is missing from GaussTR’s output but is present in EasyOcc’s. GaussTR excels at segmenting dynamic objects, producing more complete representations, while EasyOcc only predicts the visible portions, leading to hollow or partial reconstructions. This aligns with the quantitative results in Table III, where GaussTR performs better in dynamic object segmentation. However, EasyOcc benefits from a voxel-based scene representation, enabling smooth predictions, while GaussTR’s Gaussian-based representation produces more fragmented outputs.
The use of our 3D pseudo-labels offers several advantages. Firstly, our models correctly identify the region beneath the ego vehicle as road due to temporal information during 3D pseudo-label generation. Models without our labels often fail here, as this region is typically absent in camera views. Secondly, it aids in accurately predicting the overhang of a building, highlighted in the back-right camera view, which is missing in the ground-truth labels and other models’ predictions. Finally, it is particularly effective in preventing object duplication and reducing scene densification, which significantly enhances performance during evaluation without the camera mask.
@@ V-F2 Image View Analysis
Figure 9: Image View Qualitative Analysis: Examination of semantic voxel quality across the six camera views for GaussianOcc and GaussianOcc trained with 3D pseudo-labels. Conducted on nuScenes scene-783, sample token e67f3e81225f426f8e1743af45487762. Boxes are overlaid on the RGB images to highlight specific areas for qualitative analysis.
@@
Model predictions and our 3D pseudo-labels are generated solely from camera views. Hence, we shall examine the semantic voxel predictions of GaussianOcc and GaussianOcc trained with 3D pseudo-labels from the perspective of the six surrounding cameras, as shown in Figure 9.
@@
Starting with the front left camera, we observe a mislabeling in the ground-truth data, where a pole is annotated incorrectly. GaussianOcc predicts the pole but is penalized due to the incorrect label, while our model predicts vegetation in its place. In the front view, our model mistakenly predicts a bus (yellow) and truck (purple) in the distance, likely confused by a distant car in the RGB image, but both models perform well otherwise. In the front right view, both models detect a pedestrian, but the position is inaccurate due to partial occlusion by a pole. Neither model correctly labels the fire hydrant or trash can as manmade, though both correctly identify the motorcycle.
In the back right view, our model correctly identifies the overhang of a building, which GaussianOcc fails to. Both models detect the pedestrians and a traffic cone, which is absent from the ground-truth annotation. In the back view, GaussianOcc fails to predict the road in the lower part of the image, while our model produces a continuous road surface. Additionally, GaussianOcc erroneously predicts a wall behind the bus and vehicle, an object not present in the ground-truth labels or the RGB image, and absent in our model’s output. In the back left view, our model correctly identifies a traffic light, whereas GaussianOcc only partially detects it and produces several incorrect labels for objects beyond the visible vehicles, including barrier, pedestrian, and construction vehicle, none of which are supported by the ground-truth or RGB image.
V-G Discussion
This section provides a summary of the results and analyses presented in this paper.
3D pseudo-labels demonstrated the significant benefits of integration into existing models, especially in relation to the mIoU metric when evaluating with the camera mask. For instance, OccNeRF showed a 45% mIoU improvement due to high-quality supervision from Grounded-SAM and Metric3Dv2. Critical semantic classes, such as bicycle, car, and pedestrian, saw substantial gains, with the IoU for pedestrian increasing by over 627%. Our standalone model, EasyOcc, achieved a respectable performance with an mIoU of 13.86. While these results are promising, they remain behind the Gaussian-based models, GaussianFlowOcc and AutoOcc, which benefit from the incorporation of temporal information and Vision-Language Models (VLMs).
When evaluating the models without the camera mask, it becomes evident that the 3D pseudo-labels contribute to a more holistic representation of the scene, as reflected in improvements to IoU, mIoU, and per-class IoU. Notable gains were observed for SelfOcc, OccNeRF, and GaussianOcc, largely due to the reduction in object duplication and overprediction. However, this trend extends beyond these models; our approach also outperforms GaussTR, which employs a Gaussian-based scene representation. In this case, the improvements are likely due to our model producing more confident predictions with fewer incorrect outputs.
In the ablation study, we evaluated architectural and training design choices. We found only a marginal performance drop with a ResNet-18 backbone compared to ResNet-101, with an mIoU of 12.73 and a slight increase in inference speed (5.4 FPS to 6 FPS), highlighting the compensatory power of 3D pseudo-labels. The importance of aggregating temporal samples was also validated, with 10 and 13 temporal samples yielding the best results, suggesting a saturation point beyond which additional samples provide minimal gains. Further ablations on $\lambda$ in the pseudo-loss, selective removal of loss components, and refinements to Lovász-Softmax loss provided deeper insights into EasyOcc’s learning dynamics.
Qualitative analysis reinforced findings of the main results. 3D pseudo-labels produced cleaner, less noisy semantic voxel outputs and made more accurate predictions for occluded or partially visible objects, benefiting from temporal information in label generation.
Some limitations include increased memory usage and spatial resolution constraints when utilizing a voxel representation, which is necessary for our 3D pseudo-labels. To achieve finer voxel-level predictions, new 3D pseudo-labels would need to be generated and retrained at the desired resolution. This contrasts with models using discrete 3D representations, such as Gaussians, which offer more flexible resolution scaling [ttgaussocc].
Overall, we emphasize the importance of performing loss computation directly in 3D space for self-supervised models, especially when combining it with temporal aggregation from prior samples. Importantly, aggregation is only required during training, ensuring efficient deployment. By omitting 2D loss computation, we eliminate computationally expensive rendering operations, reducing training time.
@@ VI Conclusion
This paper details the use of 3D pseudo-labels for fully self-supervised semantic occupancy prediction, enabling loss computation directly in 3D space, as opposed to the conventional 2D camera space approach. Our method is easily integrable into existing architectures, yielding increased model performance and more holistic scene representation. The effectiveness of these 3D pseudo-labels highlights their potential to significantly enhance self-supervised models. However, several avenues for future work remain to further refine and evaluate the proposed approach:
@@
Self-supervised semantic occupancy prediction models have historically lagged behind supervised counterparts. However, recent advancements, including those presented in this work, suggest that self-supervised methods are rapidly gaining traction and may soon close the performance gap. While research trends shift toward discrete Gaussian representations, the widespread adoption of 3D pseudo-labels for voxel scene representation models remains uncertain. Nonetheless, this study shows that incorporating temporal information and carefully considering the domain for loss computation are crucial for achieving optimal performance in semantic occupancy prediction.

### v4

@@ V-E5 LOVÀSZ SOFTMAX LOSS
Following the previous experiment, we retrain the model with the Lovász-Softmax loss that excludes the empty index from the loss computation, seen in Table IX.
TABLE IX: Lovász-Softmax empty index ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
@@
14.9
15.4
Excluding the empty class from the loss computation results in a 0.5 mIoU improvement. Only the previous experiments in the ablation section were conducted with the inclusion of the empty label in the Lovász loss. However, the integrity of the comparative results is maintained, as all models were subject to the same conditions and limitations.
@@ V-F QUALITATIVE RESULTS
In this section, we present a qualitative analysis in two parts. First, we compare predicted semantic voxels against ground-truth labels. Then, we compare voxel estimations across the six camera views for GaussianOcc and its variant trained with our pseudo-loss to assess the benefits of Easy3D-Labels.
@@ V-F1 VOXEL ANALYSIS
Figure 7: Voxel qualitative analysis of models evaluated in Table III on sample token e67f3e81225f426f8e1743af45487762. +(Ours) indicates the same model trained with Easy3D-Labels.
We begin by analysing Figure 7, which shows semantic voxel visualisations for each model, comparing original models with their pseudo-loss variants.
SelfOcc performs well in road segmentation but misses vehicles in the back camera view; pseudo-loss corrects this but introduces increased misclassification as barrier, particularly in the back-left view, likely due to confusing structures such as fences or signs.
OccNeRF predicts more correct objects than SelfOcc, including vehicles in the back view, but suffers from object duplication, leading to penalties in metrics. Pseudo-loss reduces duplication and improves vegetation estimation. Similar improvements are observed in GaussianOcc, where pseudo-loss corrects misclassifications such as road being predicted as a wall.
Comparing EasyOcc and GaussTR, GaussTR fails to capture vegetation and building overhangs, while EasyOcc predicts these correctly. However, GaussTR better reconstructs dynamic objects, producing more complete shapes, whereas EasyOcc predicts only visible regions. This aligns with quantitative results, where GaussTR performs better on dynamic classes. EasyOcc, using a voxel representation, produces smoother outputs, while GaussTR’s Gaussian representation appears more fragmented.
Easy3D-Labels provide several benefits: they enable correct estimations of regions beneath the ego vehicle through temporal aggregation, improve detection of structures such as building overhangs absent in ground truth, and reduce object duplication and scene densification, leading to improved mIoU and RayIoU.
@@ V-F2 IMAGE VIEW ANALYSIS
Figure 8: Image view qualitative analysis of models evaluated in Table III on sample token e67f3e81225f426f8e1743af45487762. +(Ours) indicates the same model trained with Easy3D-Labels.
@@
Model estimations and our 3D pseudo-labels are generated solely from camera views. We therefore examine semantic voxel estimations of GaussianOcc and its variant trained with Easy3D-Labels across the six camera views in Figure 8.
@@
In the front left view, a ground-truth mislabel causes GaussianOcc to be penalised despite correctly predicting a pole, while our model predicts vegetation. In the front right view, both models detect a pedestrian but with inaccurate positioning due to occlusion; neither correctly labels smaller objects such as the fire hydrant or trash can, though both detect the motorcycle.
In the back right view, our model correctly captures a building overhang, which GaussianOcc misses. In the back view, GaussianOcc fails to predict the road in the lower region and incorrectly introduces a wall behind vehicles, while our model produces a continuous road and avoids this error. In the back left view, GaussianOcc generates several unsupported estimations (e.g., barrier, pedestrian, construction vehicle), whereas our model produces cleaner and more consistent outputs.
@@ VI CONCLUSION
This paper presents the use of 3D pseudo-labels, Easy3D-Labels, for self-supervised semantic occupancy estimation models for automated vehicle perception. These labels enable loss computation directly in 3D space, rather than relying on conventional 2D camera-space supervision. They can be easily integrated into existing architectures, leading to improved model performance, more complete scene representation, and better detection of vulnerable road users. Additionally, using only these labels for supervision in our model, EasyOcc, proves effective across performance metrics. The strong performance of these 3D pseudo-labels highlights their potential to enhance self-supervised models. However, several directions for future work remain to further refine and evaluate the proposed approach:
@@
Self-supervised occupancy estimation models have historically lagged behind supervised methods, but recent advancements, including this work, indicate they are closing the performance gap. While research is shifting toward Gaussian representations, the adoption of 3D pseudo-labels for voxel-based models remains uncertain. Nonetheless, this study shows that leveraging temporal information and carefully selecting the loss domain are key to achieving strong performance in semantic occupancy estimation.
Acknowledgement
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.